# Osmosis: RGBD Diffusion Prior for Underwater Image Restoration

### v1

@@ Abstract
Underwater image restoration is a challenging task because of strong water effects that increase dramatically with distance. This is worsened by lack of ground truth data of clean scenes without water. Diffusion priors have emerged as strong image restoration priors. However, they are often trained with a dataset of the desired restored output, which is not available in our case. To overcome this critical issue, we show how to leverage in-air images to train diffusion priors for underwater restoration. We also observe that only color data is insufficient, and augment the prior with a depth channel. We train an unconditional diffusion model prior on the joint space of color and depth, using standard RGBD datasets of natural outdoor scenes in air. Using this prior together with a novel guidance method based on the underwater image formation model, we generate posterior samples of clean images, removing the water effects. Even though our prior did not see any underwater images during training, our method outperforms state-of-the-art baselines for image restoration on very challenging scenes. Data, models and code are published in the project page osmosis-diffusion.github.io.
@@
Diffusion Models Underwater Image Restoration
@@ 1 Introduction
Figure 2: The iterative sampling process starts in $t=T$ with random noise in $4$ channels. The denoising step outputs $\mu$ and denoised samples $\hat{x}_{0}=(\hat{J}_{0},\hat{D}_{0})$. We use the physical image formation model together with $\hat{x}_{0}$ to optimize the water parameters $\hat{\phi}$. These are then input to the guided sampling. This process repeats itself, gradually updating both the estimated image and depth, until $t=0$, in which $x_{0}$ holds the method’s estimate for both the reconstructed scene $J_{0}$ and its depth $D_{0}$.
@@
We propose to train a prior model of RGBD images using available datasets of natural outdoor scenes that were collected in air. Using in-air scenes for underwater images may be counter-intuitive, but actually has strong benefits: 1) overcomes the lack of clean underwater image data; 2) leads to a strong prior that captures the joint statistics governing color and depth in natural scenes, where, as opposed to underwater images, color does not fade with distance; 3) prevents overfitting to specific types of underwater images.
@@ 2.1 Underwater Image Restoration and Depth Estimation
Classic image restoration. To cope with the ill-posed nature of the problem, earlier works on underwater image restoration introduced tailored image priors. Some priors aim to estimate a depth or transmission map of the scene to reduce the number of the unknowns and then use the estimated depth to restore the scenes [42, 41, 8].
@@
CWR [25] introduced the HICRD dataset, where the images are restored using optical parameters measured using ocean optics instruments. This is an impressive effort. Unfortunately, the images are acquired in a downward-looking position, and thus their depth range is very limited. In FUnIE-GAN [28] a training dataset is generated by having humans select good images from a large set of unprocessed underwater images. The humans are instructed to choose images where the foreground objects are identifiable. These are then distorted by a GAN to produce the paired poor images. A synthetic dataset was generated in [34]. The dataset is synthesized from the NYU-v2 RGBD dataset [46], using the image formation model equation and several sets of values for the water parameters.
@@ 2.2 Diffusion Model Prior
Diffusion models. Diffusion models have emerged as a powerful type of a generative model, especially for images. In the last few years several formulations and variations have been developed [47, 50, 26, 52], most of which use a U-Net architecture [43] as a noise predictor. Because the training relies on very large datasets and is extremely time-consuming, significant work has been devoted to the setup where models are first pre-trained on large datasets and only later fine-tuned to more specific data, closer to the tasks at hand [44, 48]. In DepthGen [45] this approach is taken forward, by using a pre-trained model to kickstart a new model trained on different modalities using a different architecture. This is done by replacing the input and output layers of the pre-trained U-Net. We take a similar approach for training our RGBD prior, by using a pretrained diffusion model that was trained on RGB only.
@@
Diffusion model as a prior for clean images. In addition to conditional generation of images, there is a growing body of research where diffusion models are being used as clean image priors, and image restoration is formulated as posterior sampling  [10, 13, 14, 15, 22, 51, 19, 12, 29, 30, 49, 6]. These include tasks like denoising, inpatining, deblurring, and more general tasks. The image degradation model is used to form the likelihood term and sampling from the posterior is implemented by adding the gradient of the degradation model in each step of the inverse sampling iterations, a method also known as guidance. This can be implemented in different ways, differing by the way the likelihood term is approximated, and whether its computation and gradient are based on an intermediary noisy image or on the estimated clean image in each sampling step.
The limited access to clean data in many cases, has led to research on training a prior of clean images, using noisy training data only. In [1, 31, 16] a diffusion model prior is trained using noisy data, by assuming a known degradation model. Since this setup is not directly applicable in our case, both because of the model structure and the unknown parameters, we chose instead to train our prior on clean data that was not taken underwater.
@@ 3.2 Diffusion Models
Diffusion models have proven to be very effective in capturing the distribution of natural images, given in training data. We describe here the formulation that we use in our work, adopted from [17].
A diffusion model is defined by a Markov chain designed to transform the distribution of real images $x_{0}$ to a Gaussian distribution $x_{T}$, by gradually scaling down the image values and adding Gaussian noise, using a schedule determined by the scalar parameters $\alpha_{t}$,
@@
where $\epsilon_{\theta}(x_{t},t)$ is the neural network trained to approximate $\epsilon$, and $\bar{\alpha}_{t}=\prod_{s=0}^{t}\alpha_{t}$. The same neural network can also be used to predict a diagonal covariance $\Sigma_{\theta}(x_{t},t)$. At any iteration, an estimate of the clean image $x_{0}$ can be derived by computing the mean of $p(x_{0}\mid x_{t})$:
@@
Score-based diffusion models represent another view of the above formulation, where the inverse sampling process is shown to be equivalent to an iterative stochastic differential equation (SDE) solver, and the forward process is assumed to be a continuous time differential equation. In that case it can be shown that $\epsilon_{\theta}(x_{t},t)$ is an approximation to the score function $\nabla_{x_{t}}\log p_{t}(x_{t})$.
@@ 3.3 Posterior Sampling
Under the score-based view, given an observation $y$ and a likelihood function $p(y|x)$, we can use the same mechanism to sample from the posterior $p(x|y)$, using the posterior score function.
$\nabla_{x_{t}}\log p_{t}(x_{t}|y)=\nabla_{x_{t}}\log p_{t}(x_{t})+\nabla_{x_{t}}\log p_{t}(y|x_{t})$ (6)
@@
The various works on posterior sampling propose different approximations of this integral. Some propose to collapse the uncertainty in $y$ [10, 13, 14, 15, 22, 51]. In [19] a variational inference approximation is proposed, and in [12, 29, 30, 49, 6] the likelihood model is computed on either $x_{t}$ directly, or on the mean $\hat{x}_{0}=\mathbb{E}[x_{0}|x_{t}]$, as given in Eq. 5. After some experimentation, we decide to use the latter approximation as formulated in DPS [12]:
@@ 4 Method
4.1 Our Approach
The approach in this work is to use the underwater image formation model (Eq. 1) as a forward model that maps the space of natural images $x$ to the space of underwater observed images $y$. We aim to use this forward model to construct a likelihood term $p(y\mid x)$, and use it to sample from the posterior distribution of images (Eq. 7).
In our case this cannot be implemented directly since the image formation model contains unknown parameters, namely the depth $D$ at each pixel, and the water parameters, ${\phi_{a}},{\phi_{b}},{\phi}^{\infty}$, which can differ between scenes. One way to deal with the unknown depth, is to use a monocular depth estimator in a separate first stage, and then use it as part of the forward model. In the results section we discuss this approach and show that it is suboptimal (method variant termed DA-Osmosis).
@@ 4.1 Our Approach
where $\bm{\phi}=({\phi_{a}},{\phi_{b}},{\phi}^{\infty})$ are the remaining 9 unknown water parameters. We optimize $\bm{\phi}$ during sampling, in a similar way to GDP [18], using gradient descent on the image reconstruction loss. The likelihood term used for guidance is defined as a Gaussian around $f_{\bm{\phi}}(\hat{x}_{0})$ with a fixed variance.
@@ 4.3 Sampling from the Posterior
Running posterior sampling, guided by a reconstruction loss only, we observe two problems. First, for pixels with large depth values, the reconstructed image is dominated by the backscatter, making the estimation of the clean pixel color values unstable. Second, the color values can shift outside the valid range, causing color saturation.
To overcome the first issue we multiply the reconstruction loss in every pixel by the estimated depth value (without passing the gradients of this operation, denoted by sg):
@@
where $\hat{J}_{0}(i,c)$ is the value of color channel $c$ of pixel $i$, assuming the valid color range is $[-1,1]$, $T_{v}$ and $T_{a}$ are thresholds set to values close to 1 and 0 respectively, and $\lambda_{v},\lambda_{a}$ are the scalar weights of both losses. The total loss is then given by:
@@ 5 Results
5.1 Real-World Scenes
@@
Figure 6: Example results on UIEB [35], where our method yields better results than the dataset’s ground truth. See for example the color of the sand, the orange color of the fish in the top-left, and the divers’ authentic skin color in the bottom-right.
5.2 Simulation
We conducted a simulation following [27, 34] using all the 1449 images from the NYU-v2 RGBD dataset [46] with randomly varying water parameters and generated the underwater images with the image formation model in Eq. 1. For fairness in comparison with other methods we applied ${\phi_{a}}={\phi_{b}}$. The quantitative results are summarized in Table 1. Our method substantially outperforms other methods in the image restoration metrics of PSNR, SSIM, and LPIPS. We emphasize that our prior was not trained on this data, or indoor data at all. Moreover, two of the methods [27, 34] we outperform were trained on this data.
@@ 5.2 Simulation
5.3 Ablation
@@
To demonstrate the effect of all parts of our methods, we conduct an ablation study of the following variants: 1. $\mathcal{L}=\mathcal{L}_{\rm rec}$ (instead of Eq. 12); 2. Removing weighting by $\hat{D}_{0}$ in Eq. 9; 3. Same guidance scale $s$ for all channels; 4. ${\phi_{a}}={\phi_{b}}$; 5. $\mathcal{L}=\mathcal{L}_{\rm rec}+\mathcal{L}_{\rm val}$ (remove $\mathcal{L}_{\rm avrg}$ from Eq. 12); 6. $\mathcal{L}=\mathcal{L}_{\rm rec}+\mathcal{L}_{\rm avrg}$ (remove $\mathcal{L}_{\rm val}$ from Eq. 12).
Numerical results on the simulation for variants #1-#3 are presented in Table 1. Ablations #4-#6 are shown only on real world images since in the simulation we do not use $\mathcal{L}_{\rm avrg}$, and ${\phi_{a}}={\phi_{b}}$. Fig. 7 presents the results of all variants on one of the scenes. We see that the additional losses are important to prevent color saturation and shift. Increasing the loss weight with depth improves restoration in further areas. Separating guidance scales between the RGB and depth channels improves depth reconstruction. Setting ${\phi_{a}}\neq{\phi_{b}}$ extends the range of the restoration.
@@ 6 Discussion
Like all diffusion models with U-Nets, our method is limited by a fixed resolution and long running time. This could potentially be improved using other network architectures and sampling methods. In the future, we plan to explore this direction and additional tasks that can exploit the RGBD prior.

### v2

# osmosis-diffusion.github.io

@@ Abstract
Underwater image restoration is a challenging task because of water effects that increase dramatically with distance. This is worsened by lack of ground truth data of clean scenes without water. Diffusion priors have emerged as strong image restoration priors. However, they are often trained with a dataset of the desired restored output, which is not available in our case. We also observe that using only color data is insufficient, and therefore augment the prior with a depth channel. We train an unconditional diffusion model prior on the joint space of color and depth, using standard RGBD datasets of natural outdoor scenes in air. Using this prior together with a novel guidance method based on the underwater image formation model, we generate posterior samples of clean images, removing the water effects. Even though our prior did not see any underwater images during training, our method outperforms state-of-the-art baselines for image restoration on very challenging scenes. Our code, models and data are available on the project’s website.
@@
Diffusion Models Physics-Based Computer Vision Underwater Image Restoration
@@ 1 Introduction
Figure 2: The iterative sampling process starts in $t=T$ with random noise in $4$ channels. The denoising step outputs denoised samples $\hat{x}_{0}=(\hat{J}_{0},\hat{D}_{0})$. We use the underwater physical image formation model together with $\hat{x}_{0}$ to optimize the water parameters $\hat{\phi}$, and to guide the sampling towards the observed image. This process repeats itself, gradually updating both the estimated image and depth, until $t=0$, in which $x_{0}$ holds the method’s estimate for both the reconstructed scene $J_{0}$ and its depth $D_{0}$.
@@
We propose to train a prior model of RGBD images using available datasets of natural outdoor scenes that were collected in air. Using in-air scenes for underwater images may be counter-intuitive, but actually has strong benefits: 1) it overcomes the lack of clean underwater image data; 2) it leads to a strong prior that captures the joint statistics governing color and depth in natural scenes, where, as opposed to underwater images, color does not fade with distance; 3) it prevents overfitting to specific types of underwater images.
@@ 2.1 Underwater Image Restoration and Depth Estimation
Classic underwater image restoration. To cope with the ill-posed nature of the problem, earlier works introduced tailored image priors. Some priors aim to estimate a depth or transmission map of the scene to reduce the number of the unknowns and then use the estimated depth to restore the scenes [42, 41, 7].
@@
CWR [24] introduced the HICRD dataset, where the images are restored using optical parameters impressively measured using ocean optics instruments. Unfortunately, the images are acquired in a downward-looking position, and thus their depth range is very limited. In FUnIE-GAN [27] a training dataset is generated by having humans select good images from a large set of unprocessed underwater images. The humans are instructed to choose images where the foreground objects are identifiable. These are then distorted by a GAN to produce the paired poor images. A synthetic dataset was generated in [34]. The dataset is synthesized from the NYU-v2 RGBD dataset [46], using the image formation model equation and several sets of values for the water parameters.
@@ 2.2 Diffusion Model Prior
Diffusion models. Diffusion models have emerged as a powerful type of generative models. In the last few years several formulations and variations have been developed [47, 50, 25, 52], most of which use a U-Net architecture [43] as a noise predictor. Because the training relies on very large datasets and is extremely time-consuming, significant work has been devoted to the setup where models are first pre-trained on large datasets and only later fine-tuned to more specific data, closer to the tasks at hand [44, 48]. In DepthGen [45] this approach is taken forward, by using a pre-trained model to kickstart a new model trained on different modalities using a different architecture. This is done by replacing the input and output layers of the pre-trained U-Net. We take a similar approach for training our RGBD prior, by using a pretrained diffusion model that was trained on RGB only.
@@
Diffusion model as a prior for clean images. In addition to conditional generation of images, there is a growing body of research where diffusion models are being used as clean image priors, and image restoration is formulated as posterior sampling  [9, 12, 13, 14, 21, 51, 18, 11, 28, 29, 49, 5]. These include tasks like denoising, inpainting, deblurring, and more general tasks. The limited access to clean data in many cases, has led to research on training a prior of clean images, using noisy training data only. In [1, 30, 15] a diffusion model prior is trained using noisy data, assuming a known degradation model. Since this setup is not directly applicable in our case, we chose instead to train our prior on clean data that was not taken underwater.
@@ 3.2 Diffusion Models
Diffusion models have proven to be very effective in capturing the distribution of natural images, given in training data. We describe here the formulation that we use in our work, adopted from [16]. A diffusion model is defined by a Markov chain designed to transform the distribution of real images $x_{0}$ to a Gaussian distribution $x_{T}$, by gradually scaling down the image values and adding Gaussian noise, using a schedule determined by the scalar parameters $\alpha_{t}$,
@@
where $\epsilon_{\theta}(x_{t},t)$ is the neural network trained to approximate $\epsilon$, and $\bar{\alpha}_{t}=\prod_{s=0}^{t}\alpha_{s}$. The same neural network can also be used to predict a diagonal covariance $\Sigma_{\theta}(x_{t},t)$. At any iteration, an estimate of the clean image $x_{0}$ can be derived by computing the mean of $p(x_{0}\mid x_{t})$:
@@
The above formulation can also be developed from a score-based modeling approach, where it can be shown that $\epsilon_{\theta}(x_{t},t)$ is an approximation to the score function $\nabla_{x_{t}}\log p_{t}(x_{t})$.
@@ 3.3 Posterior Sampling
Under the score-based view, given an observation $y$ and a likelihood function $p(y|x)$, we can use the same mechanism to sample from the posterior $p(x|y)$, using the posterior score function
$\nabla_{x_{t}}\log p_{t}(x_{t}|y)=\nabla_{x_{t}}\log p_{t}(x_{t})+\nabla_{x_{t}}\log p_{t}(y|x_{t})\;\;.$ (6)
@@
The various works on posterior sampling propose different approximations of this integral. Some propose to collapse the uncertainty in $y$ [9, 12, 13, 14, 21, 51]. In [18] a variational inference approximation is proposed, and in [11, 28, 29, 49, 5] the likelihood model is computed on either $x_{t}$ directly, or on the mean $\hat{x}_{0}=\mathbb{E}[x_{0}|x_{t}]$, as given in Eq. 5. After some experimentation, we decided to use the latter approximation as formulated in DPS [11]:
@@ 4 Method
We use the underwater image formation model (Eq. 1) as a forward model that maps the space of natural images $x$ to the space of underwater observed images $y$. We aim to use this forward model to construct a likelihood term $p(y\mid x)$, and use it to sample from the posterior distribution of images (Eq. 7). In our case this cannot be implemented directly since the image formation model contains unknown parameters, the depth $D$ at each pixel, and the water parameters, ${\phi_{a}},{\phi_{b}},{\phi}^{\infty}$, which can differ between scenes. One way to deal with the unknown depth, is to use a monocular depth estimator in a separate first stage, and then use it as part of the forward model. In the results we discuss this approach and show that it is suboptimal (method variant termed DA-Osmosis).
@@
where $\bm{\phi}=({\phi_{a}},{\phi_{b}},{\phi}^{\infty})$ are the remaining 9 unknown water parameters (a parameter per color channel for each of $\bm{\phi}=({\phi_{a}},{\phi_{b}},{\phi}^{\infty})$). We optimize $\bm{\phi}$ during sampling, in a similar way to GDP [17], using gradient descent on the image reconstruction loss. The likelihood term used for guidance is defined as a Gaussian around $f_{\bm{\phi}}(\hat{x}_{0})$ with a fixed variance.
@@ 4.2 Sampling from the Posterior
Running posterior sampling, guided by a reconstruction loss only, we observe two problems. First, for pixels with large depth values, the reconstructed image is dominated by the backscatter, making the estimation of the clean pixel color values unstable. Second, the color values can shift outside the valid range, causing color saturation. To overcome the first issue we multiply the reconstruction loss in every pixel by the estimated depth value (without passing the gradients of this operation, denoted by sg):
@@
where $\hat{J}_{0}(i,c)$ is the value of color channel $c$ of pixel $i$, assuming the valid color range is $[-1,1]$, $T_{v}$ and $T_{a}$ are thresholds set to values close to 1 and 0 respectively, and $\lambda_{v},\lambda_{a}$ are the scalar weights of both losses. The total loss is then:
@@ 5 Results
5.0.1 Real-World Scenes.
@@
In addition, we conducted a non-reference quantitative comparison using the MUSIQ [31] measure on $50$ real-world images presented in the paper and appendix. Our method achieved the highest score. The quantitative results are summarized in Table 2.
@@
Figure 6: Example results on UIEB [35], a dataset comprised from non-linear images with ground truth produced by different enhancement algorithms. Our method yields better results than the dataset’s ground truth. See for example the color of the sand, the orange color of the fish in the top-left, and the divers’ authentic skin color in the bottom-right.
5.0.2 Simulation.
We conducted a simulation following [26, 34] using all the 1449 images from the NYU-v2 RGBD dataset [46] with randomly varying water parameters and generated the underwater images with the image formation model in Eq. 1. For fairness in comparison with other methods we applied ${\phi_{a}}={\phi_{b}}$.
The quantitative results are summarized in Table 2. Our method substantially outperforms other methods in the image restoration metrics of PSNR, SSIM, and LPIPS. We emphasize that our prior was not trained on this data, or indoor data at all. Moreover, two of the methods [26, 34] we outperform were trained on this data.
[…]
input 51.59
contrast stretch 53.75
CWR [24] 39.63
DM [53] 54.64
FUnIE-GAN [27] 42.70
GDCP [41] 56.25
IBLA [42] 54.42
MMLE [60] 56.26
semi-UIR [26] 54.99
Ucolor [33] 47.73
USUIR [19] 52.44
UW-Net [23] 46.02
waternet [35] 53.64
unveiling [6] 50.83
DA-osmosis [58] 51.23
osmosis (ours) 56.62
Table 1: Non-reference quantitative comparison on $50$ real-world images using MUSIQ [31]. Our method achieves the highest score.
@@
5.0.3 Ablation.
To demonstrate the effect of different components in our methods, we conduct an ablation study of the following variants: 1. $\mathcal{L}=\mathcal{L}_{\rm rec}$ (instead of Eq. 10); 2. Removing weighting by $\hat{D}_{0}$ in Eq. 9 3. Same guidance scale $s$ for all channels; 4. ${\phi_{a}}={\phi_{b}}$; 5. $\mathcal{L}=\mathcal{L}_{\rm rec}+\mathcal{L}_{\rm val}$ (remove $\mathcal{L}_{\rm avrg}$ from Eq. 10); 6. $\mathcal{L}=\mathcal{L}_{\rm rec}+\mathcal{L}_{\rm avrg}$ (remove $\mathcal{L}_{\rm val}$ from Eq. 10). Numerical results on the simulation for variants #1-#3 are presented in Table 2. Ablations #4-#6 are shown only on real world images since in the simulation we do not use $\mathcal{L}_{\rm avrg}$, and ${\phi_{a}}={\phi_{b}}$. Fig. 7 presents the results of all variants on one of the scenes. We see that the additional losses are important to prevent color saturation and shift. Increasing the loss weight with depth improves restoration in further areas. Separating guidance scales between the RGB and depth channels improves depth reconstruction. Setting ${\phi_{a}}\neq{\phi_{b}}$ extends the range of the restoration.
@@ 6 Discussion
Like all diffusion models with U-Nets, our method is limited by a fixed resolution and long running time. This could potentially be improved using other network architectures and sampling methods.
Acknowledgements. The research was funded by Israel Science Foundation grant #$\#1951/23$, Israeli Ministry of Science and Technology grants #$\#1001577600$$\#1001593851$, EU Horizon 2020 research and innovation programme GA $101094924$ (ANERIS), the Leona M. and Harry B. Helmsley Charitable Trust, and the Maurice Hatter Foundation. We thank Dr. Derya Akkaynak and Dr. Matan Yuval for substantial data contribution, Amir Dayan for the paper’s name and Meirav Keidar for graphics design.
@@ Appendix 0.A Data Preprocessing and Training
Details on the training process. We train both $\epsilon_{\theta}(x_{t},~{}t)$ and $\Sigma_{\theta}(x_{t},~{}t)$ (defined in Sec. 3.2 in the paper). We use the two losses suggested in [16], $L_{\rm simple}$ is MSE with a mask for the non-valid pixels (e.g., holes and horizon) in the image, and $L_{\rm vlb}$, a Variational Lower Bound.